OpenStudy (anonymous):

Given the formula, solve for the indicated variable. f=9/5 c+32 for c

OpenStudy (anonymous):

Can you rearrange it to make c the subject?

OpenStudy (anonymous):

don't no that's how it is

OpenStudy (anonymous):

Subtract both sides by 32: F - 32 = 9C/5 + 32 - 32 F - 32 = 9C/5 Finally, multiply both sides by 5/9 to get: 5/9 * (F - 32) = 9C/5 * 5/9 5/9(F - 32) = C I hope this helps!

OpenStudy (anonymous):

U beat me again :(

OpenStudy (anonymous):

:D